OpenStudy (anonymous):

72 − 3 x (5 − 2)2? Help pleasee!

OpenStudy (anonymous):

is that x a variable factor? or just a multiplication symbol?

OpenStudy (anonymous):

thats the part that was confusing me put i think it is just a variable because there is a parenthesis meaning you already multiply.

OpenStudy (anonymous):

Well it doesn't make much sense to be a variable if it's expected to simplify to a single number. But if it's supposed to represent a line we'd have a y= that whole thing. Do you have any more context?

OpenStudy (anonymous):

nope, but i did everything and the last two is at the top sorry. if you do not understand thank you very much i think i got it a little bit.

OpenStudy (anonymous):

it was 22. somehow. i got it wrong though i thought it was 414

OpenStudy (anonymous):

\[72-3*(5-2)^2\] ?

OpenStudy (anonymous):

\[72 - 3*(5-2)^2 = 72 - 3*3^2 = 72 - 9*3 = 72-27 = 45\]

OpenStudy (anonymous):

thanks for your help you saved my lifee. lol